OpenStudy (anonymous):

Find all solutions to the equation: sin^2 x + sin x = 0

OpenStudy (mayankdevnani):

\[\large \bf \sin^2x+sinx=0\] Take sinx common, \[\large \bf sinx(sinx+1)=0\] so either sinx=0 or (sinx+1) is 0

OpenStudy (mayankdevnani):

and do remaining part yourself .. @ic0r understood?

OpenStudy (anonymous):

Here’s what I did: sin x = i a^2 + a = 0 a(a+1) = 0 * a =0 * a = -1 sin x = 0 => x = 180˚ sin x + 1 = 0 => x = 270˚ I convert those angles to radians: π, and 3π/2. Am I doing this correctly?

OpenStudy (anonymous):

It’s the latter part that I am not sure about @mayankdevnani

OpenStudy (anonymous):

I don’t think π and 3π/2 are all the solutions

OpenStudy (mayankdevnani):

See, if \(\large \bf six=0\),then the solution is :- \[\large \bf x=n \pi ,~where~ n \in I\]

OpenStudy (mayankdevnani):

and if sinx=-1, then the solution is :- \[\large \bf x=(4n-1)\frac{\pi}{2},~where~n \in I\]

OpenStudy (mayankdevnani):

hope you understand. @ic0r

OpenStudy (mayankdevnani):

and also \[\large \bf \color{red}{Welcome}~To~\color{blue}{Open}~\color{green}{Study}\]

OpenStudy (anonymous):

You lost me at x = nπ.

OpenStudy (mayankdevnani):

see when sinx=0, REMEMBER :- general solution of x=n pi where n is integer like,x=pi,2 pi etc. all give the value ZERO. or you can say that x=180 degrees,360 degress etc.-----> Value=0

OpenStudy (anonymous):

I actually have no idea what you are talking about, I'm really confused. Is the step that I got x = pi and 3pi/2 correct?

OpenStudy (mayankdevnani):

yeah...you are correct. I am explaining that how this value come about ? By the way,good job :) @ic0r

OpenStudy (anonymous):

So those are all the solutions of the equation? I'm in doubt for a little bit.

OpenStudy (mayankdevnani):

yeah !!

OpenStudy (anonymous):

That's it? Have I been overthinking?

OpenStudy (mayankdevnani):

what ?

OpenStudy (mayankdevnani):

@ic0r there is a rule of OPEN STUDY that if anyone gave your answer correctly,then you have to give MEDAL to that user. To give medal,there is option bar BEST RESPONSE right side to my comment and press it,you gave medal to that user.

OpenStudy (mayankdevnani):

understood? @ic0r

OpenStudy (anonymous):

So I have been correct on my answer?

OpenStudy (mayankdevnani):

yeah! absolutely correct

OpenStudy (anonymous):

Thank you so much for your help! I appreciate it a lot! @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)